# Resource-constrained Fairness

### v4

@@ Abstract
Access to resources strongly constrains the decisions we make. While we might wish to offer every student a scholarship, or schedule every patient for follow-up meetings with a specialist, limited resources mean that this is not possible. When deploying machine learning systems, these resource constraints are simply enforced by varying the threshold of a classifier. However, these finite resource limitations are disregarded by most existing tools for fair machine learning, which do not allow the specification of resource limitations and do not remain fair when varying thresholds. This makes them ill-suited for real-world deployment. Our research introduces the concept of “resource-constrained fairness" and quantifies the cost of fairness within this framework. We demonstrate that the level of available resources significantly influences this cost, a factor overlooked in previous evaluations.
@@ 1 Introduction
Machine learning models are used to make decisions in high-impact areas of our lives such as finance, justice, and healthcare (Mehrabi et al., 2021). Fair machine learning has emerged in response to the notion that simply making maximally accurate decisions is not enough and that training high-performance classifiers can result in both the transfer of existing biases from data to new decisions, as well as the introduction of new biases (Wachter et al., 2020).
@@
Many studies that focus on improving fairness in machine learning overlook the practical limitations under which these models operate. For example, scenarios including university admissions, healthcare provision, and corporate hiring, are normally constrained by finite resources. Universities have a restricted quota of students to admit annually, healthcare facilities are bounded by available space and staff, and companies have a limited number of positions to fill. Even for banking, where in principle banks can keep making loans providing enough people pay them back, for any particular time period they will have a limited set of resources and only be able to loan out a certain amount. In all these cases, once the available resources are fully used, increasing the selection rate of disadvantaged groups must necessarily involve reducing selection from more privileged groups.
Figure 1: Illustration of how the different bias mitigation methods enforce fairness at a specific selection rate (or threshold) but break when we consider the full range of possible selection rates (Adult Income dataset) . [^2] The icons represent the labels of the bias mitigation methods at one specific threshold. The dashed line represent the output for the other thresholds, by using the prediction scores generated by that method. Note that this approach is not possible for Fairlearn as they do not return a prediction score.
@@
However, limited resources are not taken into account in most discussions in fair machine learning. Kwegyir-Aggrey et al. (2023) also highlight how practitioners typically need to adapt the threshold to ensure that outcomes meet their domain-specific needs, while corbett2023measure argue that the trade-offs in fair decision-making are most acute when there are constrained resources. This failure to account for real-world constraints on fairness may be a contributing factor to the field’s lack of prior cases of enforcing fairness on production models, which contrasts starkly with the abundance of studies using fairness metrics solely for model testing and debugging.
Figure 2 illustrates how demographic parity, one of the most common notions of fairness, varies with selection rate. Many of the currently popular bias mitigation methods, enforced using a range of tools from AIF360 Bellamy et al. (2018) and Fairlearn Bird et al. (2020), aim to ensure fairness at a specific threshold. However, as the thresholds vary, fairness oscillates wildly, revealing that these methods may break when considering the whole range of possible selection rates.
@@
In the standard setting of an unconstrained budget, Mittelstadt et al. (2023) posit that increasing harm to privileged groups solely to achieve fairness is not an optimal approach and refer to this as ‘leveling down’. We agree with this viewpoint but argue that within contexts constrained by limited resources, some rebalancing is necessary to redistribute the resources. In these circumstances, deliberately not fully utilizing resources to ensure equality could also be termed ‘leveling down’. .
In this study, we make the following contributions:
@@
  • We reformulate the problem of fair machine learning as a resource-constrained problem, where we treat the positive decisions as resources to allocate among different groups. We propose an algorithm that enforces fairness under resource limitations.
@@
  • We provide a connection to leveling up, demonstrating that under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
  • We quantify the cost of fairness in this resource-constrained framework. We study the impact of factors inherent to the dataset. We find that the selected positive decision rate has a substantial impact on this cost, an aspect that is not taken into account in previous evaluations.
  • We explore the practical implications of our findings and discuss what strategic actions practitioners might undertake to mitigate trade-offs between groups when fairness is constrained by resource limitations.
@@ 2 Background
We consider fairness in classification. Starting with a classifier $c_{w}(\cdot)$ parameterized by weights $w$. Let $\mathrm{H}_{\mathcal{D}}[c_{w}]$be a measure of expected harm of a classifier $c_{w}(\cdot)$ over a particular distribution $\mathcal{D}$ where $y_{x}$ represents the true target label of instance $x$, i.e.,
@@
Such measures of a harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$might be $1-\text{precision}$ (for example, when measuring the proportion of people incorrectly stopped by the police); or $1-\text{recall}$ (when measuring the number of cancer cases that go unflagged for follow-up treatment).
In fair classification, we usually want to measure fairness with respect to a protected attribute, such as gender or ethnicity. Using this protected attribute, we can partition the dataset into groups. Group fairness metrics measure (in)equality between these groups with respect to some measure of harm. For example, equal opportunity requires the recall between groups to be equal, while demographic parity enforces an equal selection rate (Verma and Rubin, 2018). Typically, a fair classifier is found by minimizing some global loss $\ell$ (such as accuracy or a continuous proxy such as the logistic loss) while ensuring that the harm is the same per group.
@@
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))\text{ such that }$ (2)
$\mathrm{H}_{g_{1}}[c_{w}]=\mathrm{H}_{g_{2}}[c_{w}]\,\forall g_{1},g_{2}\in\mathcal{G}$
Typical notions of fairness include demographic parity, where the harm corresponds to $1-$ the selection rate, and equal opportunity where it corresponds to the false negative rate.
@@ 2.1 Leveling Down
Mittelstadt et al. (2023) observe that methods to enforce fairness often level down; that is, they may enforce fairness by decreasing harm in some groups, but also by increasing harm to other groups (e.g. naively enforcing Equal Opportunity while detecting cancer will often result in a higher false negative rate for some groups i.e. they receiving a lower rate of cancer detection than they would otherwise). This process of enforcing fairness can alter the overall selection rate of the model (Goethals et al., 2024), and is distinct from any leveling down that might be inherently required by the resource constraints discussed earlier. To this end, they suggested replacing Equation LABEL:eq:fair_class with rate constraints, which enforced that the harm in question (e.g. being denied follow-up care) should be below a certain level $k$ for every group:
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))\text{ such that }\mathrm{H}_{g}[c_{w}]\leq k\,\forall g\in\mathcal{G}$ (3)
@@
This is the same as saying that in the worst case, the harm should be below k, for any group. The challenge now becomes, “How should $k$ be set?” When deciding to deploy a particular model, stakeholders and data scientists often have to agree on acceptable global levels of harm, for example an acceptable recall rate (referred to as sensitivity in the medical literature) for early cancer detection. A similar process can be used to select $k$ instances per group.
@@
This notion of leveling-up, and decreasing the harm in Equation 3 is related to but distinct from minimax-fair machine learning, which minimizes the error for the worst-off group (Martinez et al., 2020). In both minimax fairness and leveling-up the concern is with reducing the harm to the worst-off group and only increasing harms to other groups when necessary, but in minimax fairness the harm is exclusively assumed to be caused by a lack of accuracy or high log-loss (this leads to a formulation of minimize the maximal loss, hence minimax). In comparison, leveling-up is concerned with decreasing other harms such as per-group recall or selection rate, which are distinct from accuracy.  [^3] Both notions are closely aligned to the philosophical principle of the ‘maximin rule’ (Rawls, 2017). According to Rawls, resources should be distributed such that they maximize the benefits for the least advantaged members of society. However, as a fundamental limitation, neither minimax nor leveling-up considers what should be done under constrained resources.
@@
Taking a step back, we can ask the questions: “If $H$ is a harm, why not set $k$ to zero? If particular groups are being harmed by not being selected, why not select everyone, and not bother with machine learning? Similarly, if people are being harmed by a failure to detect cancer, why not schedule everyone for follow-up testing?”
@@
There are multiple possible answers here. One is a matter of personal utility, that harms are generally not one-sided. Failure to repay a loan can lead to bankruptcy for both the lender and customer, resulting in devastating personal consequences. Scheduling unnecessary medical tests is at best alarming, and in the worst case can result in death, depending on how intrusive the follow-up tests are. These types of considerations have been addressed by prior work on the cost of fairness, where personal utility is measured by accuracy, (Menon and Williamson, 2018) and fairness on the ground, which measures the effect of the actual implementation of fairness principles in real-world situations (Bakalar et al., 2021). Constrained resources provide another answer. Ideally, we would prefer to offer scholarships to every student and fast-track treatment for every patient. However, real-world scenarios often lack the necessary resources to achieve this. Understanding how fairness choices are limited and guided by resource constraints is thus crucial to move fair machine learning from a testing to production environment.
Fair classification under limited resources is also related to the domain of fair allocation within welfare economics. This field focuses on ensuring that resources such as time or physical goods are distributed amongst actors in a way that meets certain criteria. This leads to a tension between maximum utility and fairness (Bertsimas et al., 2011, 2012; Donahue and Kleinberg, 2020). Elzayn et al. (2019) describe a resource allocation problem as a problem where there is a limited supply of resources to be distributed across multiple groups with different needs, such as allocating loans or disaster response. We suggest that when the decision-maker has limited resources, all fair classification tasks can be interpreted as resource allocation problems, where positive decisions represent the resources to be distributed.
2.2 Constrained Resources
@@
We formalize the notion of resources, as the number of instances predicted as positive by the machine learning model. This term can be used interchangeably with capacity, or with selection rate, positive decision rate or positive prediction rate when speak about the proportion of positively predicted instances. We define $\mathrm{r}_{\mathcal{D}}[c_{w}]$to be the selection rate (i.e. the expected proportion of positive decisions) of a classifier over a distribution $\mathcal{D}$.
@@
Under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
@@
We consider the problem of optimizing the distribution of limited resources to minimize the maximum harm experienced by any group, given by:
$\min\max_{g}\mathrm{H}_{g}[c_{w}]\text{ such that }\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$ (4)
where $\mathrm{r}_{\mathcal{D}}[c_{w}]$is a weighted sum of group-specific rates $\mathrm{r}_{g}[c_{w}]$ with all weights positive, thus making $\mathrm{r}_{\mathcal{D}}[c_{w}]$strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$.
We first consider the case where $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing as a function of $\mathrm{r}_{g}[c_{w}]$ (for example, if the harm is $1-\text{recall}$, we expect the harm to be strictly decreasing when more individuals of the group are selected). As $\mathrm{H}_{g}[c_{w}]$ can be written as an invertible function of $\mathrm{r}_{g}[c_{w}]$, $\mathrm{r}_{g}[c_{w}]$ is also strictly decreasing with respect to $\mathrm{H}_{g}[c_{w}]$. Since $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with $\mathrm{r}_{g}[c_{w}]$, and $r_{D}$ is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$, it follows that $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with respect to $r_{D}$.
We define $W$ as the set of worst-off groups such that:
@@
The optimum must occur when $\mathrm{r}_{\mathcal{D}}[c_{w}]=r$. If not, one could distribute the remaining resources such that the rates $\mathrm{r}_{g}[c_{w}]$ for the worst-off groups increase without exceeding r, reducing the maximum harm (since $\mathrm{H}_{g}[c_{w}]$ is decreasing with respect to $\mathrm{r}_{g}[c_{w}]$).
Furthermore, the optimum is reached when the harm levels are equal across all groups (so all groups are in $W$). If this were not the case, then there would be at least one group not in $W$, that could sacrifice some of its resources to the worse-off groups. Adjusting the allocation in such a way that the group not in $W$ receives less, while groups at maximal harm receive more, would lead to a reduced overall maximum harm due to the strict decrease of $\mathrm{H}_{g}[c_{w}]$ with $\mathrm{r}_{g}[c_{w}]$. Thus, under these conditions, equality in harm distribution is enforced in the optimal solution, ensuring that the distribution of resources maximally benefits the least advantaged groups within the constraints set by $\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$.
The proof also holds when the harm considered is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$ (e.g., a harm corresponding to $1-\text{precision}$). Here, the optimal solution is trivial with an overall selection rate of 0, and equality will be preserved. However, in this case, it might be more realistic to flip the sign and require the overall selection rate to be at least some r. We can simply replace the classifier $c_{w}$ with its negation, and now find that the harm is strictly decreasing with respect to the selection rate of the new classifier. Revisiting the previous proof, we find that equality holds.
Hence, we find that equality must hold in all four cases (strictly increasing or strictly decreasing harm, and a global selection rate that is either constrained above or below some value). ∎

### v5

@@ Abstract
Access to resources strongly constrains the decisions we make. While we might wish to offer every student a scholarship, or schedule every patient for follow-up meetings with a specialist, limited resources make this infeasible. When deploying machine learning systems, these resource constraints are typically enforced by adjusting the classifier threshold. However, these finite resource limitations are disregarded by most existing tools for fair machine learning, which do not allow for the specification of resource limitations and do not remain fair when varying thresholds. This makes them ill-suited for real-world deployment. Our research introduces the concept of “resource-constrained fairness" and quantifies the cost of fairness within these constraints. We demonstrate that the level of available resources significantly influences this cost, a factor overlooked in prior evaluations.
@@ 1 Introduction
Machine learning models are used to make decisions in many high-impact areas of our lives such as finance, justice, and healthcare Mehrabi et al. (2021). Fair machine learning has emerged in response to the notion that simply making maximally accurate decisions is not enough and that training high-performance classifiers can result in both the transfer of existing biases from data to new decisions, as well as the introduction of new biases Wachter et al. (2020). Many studies that focus on improving fairness in machine learning overlook the practical limitations under which these models operate. For example, scenarios including university admissions, healthcare provision, and corporate hiring, are usually constrained by finite resources. Universities have a restricted quota of students to admit annually, healthcare facilities are bounded by available space and staff, and companies have a limited number of positions to fill. Even for banking, where in principle banks can keep making loans providing enough people pay them back, for any particular time period they will have a limited set of resources and only be able to loan out a certain amount. In all these cases, once the available resources are fully used, increasing the selection rate of disadvantaged groups must necessarily involve reducing selection from more advantaged groups.
@@
However, limited resources are not taken into account in most discussions in fair machine learning. This is illustrated in Figure 1, where for many bias mitigation methods fairness oscillates wildly over different resource levels. Kwegyir-Aggrey et al. (2023) highlight how practitioners typically need to adapt the threshold to ensure that outcomes meet their domain-specific needs, while Corbett-Davies et al. (2023) argue that the trade-offs in fair decision-making are most acute when there are constrained resources. The inability to account for real-world constraints may contribute to the tiny amount of fair algorithms publically being deployed. This contrasts starkly with the abundance of studies that use fairness metrics solely for the evaluation of deployed systems Buolamwini and Gebru (2018); Barocas et al. (2023).
@@
In the typical setting of an unconstrained budget, Mittelstadt et al. (2023) posit that increasing harm to advantaged groups solely to achieve fairness is not an optimal approach and refer to this as ‘leveling down’. We agree with this viewpoint, but argue that within contexts constrained by limited resources, rebalancing is often necessary to redistribute the resources. In these circumstances, deliberately not fully utilizing resources would be a form of ‘leveling down’. In this study, we make the following practical and theoretical contributions:
@@
  • We formulate fair machine learning as a resource-constrained problem, where we treat positive decisions as a resource to allocate among different groups.
@@
  • We provide a connection to leveling up Mittelstadt et al. (2023), showing that under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
  • We extend prior work on the cost of fairness (Corbett-Davies et al., 2017) by quantifying this cost within resource-constrained settings. Using underlying parameters, we derive mathematical bounds and empirically validate these patterns.
@@ 2 Background
We consider fairness in classification. Starting with a classifier $c_{w}(\cdot)$ parameterized by weights $w$, we let $\mathrm{H}_{\mathcal{D}}[c_{w}]$  be a measure of expected harm of a classifier $c_{w}(\cdot)$ over a particular distribution $\mathcal{D}$ where $y_{x}$ represents the true target label of instance $x$, i.e.,
@@
Such measures of a harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$ might be $1-\text{precision}$ (for example, when measuring the proportion of people incorrectly stopped by the police); or $1-\text{recall}$ (when measuring the number of cancer cases unflagged for follow-up treatment).
In fair classification, we typically measure fairness with respect to a protected attribute, such as gender or ethnicity. Using this protected attribute, we can partition the dataset into groups. Group fairness metrics measure the (in)equality between these groups with respect to some measure of harm. For example, equal opportunity requires the recall between groups to be equal, while demographic parity enforces an equal selection rate Verma and Rubin (2018). Typically, a fair classifier is found by minimizing some global loss $\ell$ (such as accuracy or a continuous proxy such as the logistic loss) while ensuring that the harm is the same per group.
@@
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))$ (2)
such that $\mathrm{H}_{g_{1}}[c_{w}]=\mathrm{H}_{g_{2}}[c_{w}]\,\forall g_{1},g_{2}\in\mathcal{G}$
Notions of fairness include demographic parity, where the harm corresponds to $1-$ the selection rate, and equal opportunity where it corresponds to the false negative rate.
Figure 1: An illustration of how demographic parity varies with selection rate. We apply a range of classical bias mitigation methods from AIF360 Bellamy et al. (2018) and observe how fairness varies with selection rate. The dots indicate default thresholds. As the thresholds vary, fairness oscillates wildly (results on the Adult dataset, details Section A.2.1). Just as standard fairness methods fail to consider the selection rate, existing analyses of the cost of fairness Corbett-Davies et al. (2017); Friedler et al. (2019); Haas (2019); von Zahn et al. (2021) also fail to take it into account.
@@ 2.1 Leveling Down
Mittelstadt et al. (2023) observe that methods to enforce fairness often level down; that is, they may enforce fairness by decreasing harm in some groups, but also by increasing harm to other groups (e.g. naively enforcing Equal Opportunity while detecting cancer will often result in a higher false negative rate for some groups – they receive a lower rate of cancer detection than they would otherwise). This process of enforcing fairness can alter the overall selection rate of the model Goethals and Calders (2024), and is distinct from any leveling down that might be inherently required by the resource constraints discussed earlier. To this end, they suggested replacing Equation LABEL:eq:fair_class with rate constraints, which enforced that the harm in question (e.g. being denied follow-up care) should be below a certain level $h$ for every group:
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))$ (3)
[…]
This is the same as saying that in the worst case, the harm should be below $h$, for any group.
@@
The challenge now becomes “How should $h$ be set?” When deploying a particular model, stakeholders and data scientists often have to agree on acceptable global levels of harm, for example, an acceptable recall rate (referred to as sensitivity in the medical literature) for early cancer detection. A similar process can also select a maximal per group harm.
@@
This notion of leveling-up, and decreasing the harm in Equation 3 is related to but distinct from minimax-fair machine learning, which minimizes the error for the worst-off group Martinez et al. (2020); Diana et al. (2021); Abernethy et al. (2020). In both minimax fairness and leveling-up the concern is with reducing the harm to the worst-off group and only increasing harms to other groups when necessary, but in minimax fairness the harm is exclusively assumed to be caused by a lack of accuracy or high log-loss (this leads to a formulation of minimize the maximal loss, hence minimax). In comparison, leveling-up is concerned with decreasing other harms such as per-group recall or selection rate, which are distinct from accuracy.  [^1] Both notions are closely aligned to the philosophical principle of the maximin rule Rawls (2017). According to Rawls, resources should be distributed in a way that maximizes the benefits for the least-advantaged members of society. However, as a fundamental limitation, neither minimax nor leveling-up considers what should be done under constrained resources.
@@
Taking a step back, we can ask the questions: “If $H$ is a harm, why not set $h$ to zero? If particular groups are being harmed by not being selected, why not select everyone and not bother with machine learning? Similarly, if people are being harmed by a failure to detect cancer, why not schedule everyone for follow-up testing?
@@
There are multiple possible answers here. One is a matter of personal utility that harms are generally not one-sided. Failure to repay a loan can lead to bankruptcy for both the lender and customer, resulting in devastating personal consequences. Scheduling unnecessary medical tests is at best alarming, and in the worst case can result in death, depending on how intrusive the follow-up tests are. These types of considerations have been addressed by prior work on the cost of fairness, where personal utility is measured by accuracy Menon and Williamson (2018), and Bakalar et al. (2021), which measures the effect of the actual implementation of fairness principles in real-world situations.
Constrained resources provide another answer. Ideally, we would prefer to offer scholarships to every student and fast-track treatment for every patient. However, the real world often lacks the resources needed to achieve this. Understanding how fairness choices are limited and guided by resource constraints is thus crucial to move fair machine learning from the theoretical to a production environment.
Fair classification under limited resources is related to the domain of fair allocation within welfare economics, which focuses on ensuring that resources such as time or physical goods are distributed among actors in a way that meets certain criteria Bertsimas et al. (2011, 2012); Donahue and Kleinberg (2020); Rambachan et al. (2020); Sinclair et al. (2022); Banerjee et al. (2022). Another related area is fair ranking, which often aims to ensure fairness with respect to a top-k selection Zehlike et al. (2017).
3 Constrained Resources
@@
We formalize the notion of resources, as the proportion of instances predicted as positive by a machine learning model. This term can be used interchangeably with capacity, or with selection rate, positive decision rate, or positive prediction rate when talking about the proportion of positively predicted instances.
@@
Under constrained resources, maximally leveling-up results in equality of harm between groups.
This means that when we are optimizing the distribution of a specified number of resources and are minimizing the harm experienced by any group (leveling up), this would lead to equality of harm between groups.
@@
We define $\mathrm{r}_{\mathcal{D}}[c_{w}]$ as the selection rate (i.e. the expected proportion of positive decisions) of a classifier over a distribution $\mathcal{D}$. We consider the problem of optimizing the distribution of limited resources, denoted by $r$, in order to minimize the maximum harm experienced by any group. This optimization problem is formulated as:
$\min\max_{g}\mathrm{H}_{g}[c_{w}]\quad\text{ such that }\quad\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$ (5)
where $\mathrm{r}_{\mathcal{D}}[c_{w}]$ is a weighted sum of group-specific selection rates $\mathrm{r}_{g}[c_{w}]$ with all weights positive, thus making $\mathrm{r}_{\mathcal{D}}[c_{w}]$ strictly increasing with respect to each $\mathrm{r}_{g}[c_{w}]$.
Case 1: Harm is strictly decreasing with respect to $\mathrm{r}_{g}[c_{w}]$
We first consider the case where $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing as a function of $\mathrm{r}_{g}[c_{w}]$ (for example, if the harm is $1-\text{recall}$, we expect the harm to be strictly decreasing when more members of the group are selected). Since $\mathrm{H}_{g}[c_{w}]$ can be written as an invertible function of $\mathrm{r}_{g}[c_{w}]$, it follows that $\mathrm{r}_{g}[c_{w}]$ is also strictly decreasing with respect to $\mathrm{H}_{g}[c_{w}]$.
Given that $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with $\mathrm{r}_{g}[c_{w}]$, and $r_{D}$ is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$, it follows that $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with respect to $r_{D}$.
We define $W$ as the set of worst-off groups, i.e., the groups experiencing the maximum harm, such that:
@@
The optimum must occur when $\mathrm{r}_{\mathcal{D}}[c_{w}]=r$. If not, one could distribute the remaining resources such that the rates $\mathrm{r}_{g}[c_{w}]$ for the worst-off groups increase without exceeding r, reducing the maximum harm (since $\mathrm{H}_{g}[c_{w}]$ is decreasing with respect to $\mathrm{r}_{g}[c_{w}]$). Therefore the optimal solution uses all the available resources.
Furthermore, the optimum is reached when the harm levels are equal across all groups (so all groups are in $W$). If this were not the case, then there exists at least one group that is not in $W$ and that could sacrifice some of its resources to the worse-off groups. Adjusting the allocation in such a way that the group not in $W$ receives less, while groups at maximal harm receive more, would lead to a reduced overall maximum harm due to the strict decrease of $\mathrm{H}_{g}[c_{w}]$ with $\mathrm{r}_{g}[c_{w}]$. Thus, under these conditions, equality in harm distribution is enforced in the optimal solution, ensuring that the distribution of resources maximally benefits the least advantaged groups within the constraints set by $\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$.
Case 2: Harm is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$
Now consider the case where harm is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$. An example of this might be a harm function like $1-\text{precision}$, where increasing the selection rate for group $g$ leads to higher harm. In this case, the optimal solution is trivial with an overall selection rate of 0, as this would minimize harm across all groups. Equality is (trivially) achieved as no group receives any resources. However, in this case, it might be more realistic to flip the sign and require the overall selection rate to be at least some r. We can simply replace the classifier $c_{w}$ with its negation, and now find that the harm is strictly decreasing with respect to the selection rate of the new classifier. Revisiting the previous proof, we find that equality holds.
Hence, equality must hold in all four cases (strictly increasing or strictly decreasing harm, and a global selection rate that is either constrained above or below some value). ∎